OpenStudy (anonymous):

Identify the conic section, describe the graph and it's lines of symmetry. Then find the domain and range of the equation. 3y^2-x^2=25

OpenStudy (triciaal):

it looks like a circle it could be an ellipse for a circle center (a,b) then (x-a)^2 + (y-b)^2 =r^2 where r is the radius for an ellipse x^2/a^2 +y^2/b^2 = 1